4.37. Compute the following integrals. (a) Sci-12 12adz exp(2) JC(0,2) Z(z -3) (a) exp(z) JC(0,4) Z(z - 3)
Answer #1

ANSWER:-

The following integrals, f(2)= -21 2 - 22 U-7.2 (2+Z) (2-3) Z = 27-2 2:2 Lies on (-1,2) hence, de todos 2-2. 2=2 r -2017 sint2 1 C[012] Z(2-3) poles ? (2. 3) = 0 2-0,3 Poll z=0 lies in c(0,27 hence it - Tez 23 (2-3) -dz - 2014 5-2/3+1 Ca Sc congue ?